OpenStudy (anonymous):

When you have an equation in point slope form, does it matter which points you use?

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) the point just has to be a point on the line :)

OpenStudy (igreen):

No, it doesn't..any point on the line will do. \(y - y_1 = m(x - x_1)\) Where \(y_1\) is the y-value of the point, \(x_1\) is the x-value of the point, and \(m\) is the slope.

OpenStudy (anonymous):

Thank you both. :)

OpenStudy (igreen):

No problem.